# Fractionalized superconductivity from Majorana glue in the Kitaev–Kondo lattice

### v1

@@ Abstract
Fractionalized phases of matter in space dimensions $d\geq 2$ attract widespread interest in condensed matter research: They are characterized by emergent gauge fields, topological order, and fractionalized excitations which constitute novel quasiparticles [1, 2]. Fractionalization is often discussed for bulk insulators, with examples including Mott-insulating spin liquids and fractional Quantum-Hall or Chern insulator states. Significantly less work has been devoted to fractionalized metals and superconductors, mainly because the presence of bulk charge carriers renders both conceptual and experimental distinctions between fractionalized and conventional phases more subtle than in the simpler insulating cases.
One class of fractionalized metals whose existence can be rigorously established theoretically is that of fractionalized Fermi liquids (FL) [3, 4]: A two-band model where the electrons in one strongly correlated band form a fractionalized spin liquid of local moments, e.g. because of frustration, and are weakly exchange-coupled to a second band of weakly correlated electrons realizes an FL phase. The FL phase is characterized by the coexistence of fractionalized spin excitations, and their associated gauge field, as well as charge carriers with charge $|e|$ and spin $1/2$ whose Fermi volume violates Luttinger’s theorem [5]. In the described setting, a stronger coupling between the two bands leads to Kondo screening of the local moments, resulting in a more conventional heavy Fermi-liquid (FL) phase. Consequently, the transition between FL and FL is an instance of a Kondo-breakdown quantum phase transition [6, 7, 3, 8, 9], accompanied by a jump in the Fermi volume but no spontaneous symmetry breaking.
The focus of this paper is on constructing a superconducting cousin of the FL phase, i.e., a fractionalized superconductor (SC). In fact, fractionalized superconductivity, i.e., superconductivity coexisting with topological order, has been hypothesized from early on to exist in underdoped cuprate superconductors [10]. More recently, it has been discussed [11] in the context of scenarios of the cuprate pseudogap representing a fractionalized Fermi liquid [12, 13, 14]. A transition between fractionalized and conventional superconductivity (SC) has also been proposed to occur in an iron-based superconductor [15].
@@
In this paper we discuss and solve a tractable model of fractionalized superconductivity. We consider the Kitaev–Kondo lattice model, Fig. 1, constructed from the celebrated honeycomb-lattice spin-liquid model of Kitaev [16] and a band of conduction electrons. This model has been shown to display a FL–FL transition which – at the mean-field level – is masked by non-trivial but non-fractionalized superconductivity at intermediate Kondo coupling [17, 18]. Here we focus on the fractionalized weak-coupling regime where we can perturbatively derive an effective model for conduction-electron interactions mediated by the Majorana fermions of the spin liquid. We solve this model using the functional renormalization group (FRG) technique [19, 20] and find stable superconductivity of either chiral $d$-wave ($p$-wave) type for ferromagnetic (antiferromagnetic) Kitaev coupling, respectively, over a range of band fillings. These pairing states coexist with the fractionalized spin liquid in the local-moment sector, rendering this SC phase a natural descendant of the model’s FL phase.
@@
in standard notation, with $S_{i}^{\alpha}$ representing the Kitaev spin at site $i$ and $s^{\alpha}_{i}=c_{i\sigma}^{\dagger}\tau^{\alpha}_{\sigma\sigma^{\prime}}c_{i\sigma^{\prime}}/2$ the conduction-electron spin operator where $\tau^{\alpha}$ the vector of Pauli matrices. The first term represents the conduction-electron kinetic energy, the second the Kitaev coupling among the spin-$1/2$ local moments, with $\langle ij\rangle_{\alpha}$ denoting an $\alpha$ bond on the lattice ($\alpha=x,y,z$). The last term represents the local Kondo coupling. A chemical potential $\mu$ is applied to the conduction electrons to control their filling, $n_{c}=\frac{1}{2N}\sum_{i\sigma}c_{i\sigma}^{\dagger}c_{i\sigma}$, where $N$ is the number of unit cells. We note that $n_{c}=1$ corresponds to the “half-filled” case where the chemical potential is at the Dirac point. We will concentrate on $n_{c}\geq 1$; phases for $n_{c}\leq 1$ are related by particle–hole symmetry.
The Kitaev model $\mathcal{H}_{K}$ alone describes an exactly solvable $\mathbb{Z}_{2}$ spin liquid [16]. Its degrees of freedom are itinerant “matter” Majorana fermions and static $\mathbb{Z}_{2}$ gauge fluxes. The matter-Majorana spectrum is gapless and of Dirac type for isotropic couplings, $K^{x}=K^{y}=K^{z}\equiv K$, but acquires a gap for large anisotropies. In this paper, we will assume isotropic Kitaev couplings as well as isotropic Kondo couplings, $J_{\rm K}^{x}=J_{\rm K}^{y}=J_{\rm K}^{z}\equiv J_{\rm K}$. Upon adding the Kondo coupling, the $\mathbb{Z}_{2}$ fluxes are no longer conserved, and the solubility is spoiled.
We quickly summarize the results of Refs. 17, 18 where the Kitaev–Kondo lattice model $\mathcal{H}$ has been studied using perturbative and mean-field techniques. The model hosts a $\mathbb{Z}_{2}$ fractionalized Fermi liquid [3] for $J_{\rm K}\ll K,t$ because the Kitaev spin liquid is stable against a small coupling to conduction electrons. Conversely, the model realizes a conventional (i.e., non-fractionalized) heavy Fermi liquid for $K\ll J_{\rm K}\sim t$ (or $K\ll T_{\rm K}$ where $T_{\rm K}$ is the Kondo temperature) due to robust Kondo screening of the local moments. According to the mean-field results, the transition between these two symmetric metallic phases is masked by superconductivity, in accordance with the general considerations of Ref. 3. At the mean-field level, superconductivity sets in concomitantly with Kondo screening, hence the superconducting phase is non-fractionalized (but may have non-trivial topological band invariants). The precise character of the superconducting phases depends on the mean-field scheme and differs between Refs. 17 and 18. We note that the Kitaev–Kondo lattice has also been utilized to describe heterostructures composed of graphene and $\alpha$-RuCl3 [21].
@@
Figure 2: FRG phase diagrams of the (a) AFM and (b) FM Hubbard–Kitaev model as a function of critical scale $\Lambda_{c}$ vs. filling $n$. A magnetically ordered phase (SDW) at and near van-Hove filling $n_{\rm vH}=5/4$ is surrounded on both sides by a chiral $d$-wave, and $p$-wave superconductor, respectively. Shown are results for $U/t=2$, 2.5 and 3. The crosses indicate the absence of any divergence, i.e., the Fermi liquid prevails down to scales of $\Lambda=10^{-6}$.
@@
FRG analysis of effective model.— We analyze the leading instabilities of the Hubbard–Kitaev model by virtue of the truncated-unity FRG (TUFRG) [19, 20] method, using the recently published FRG codebase divERGe [29]. FRG interpolates between the bare Hubbard–Kitaev interaction and a low-energy two-particle interaction vertex, by means of iteratively integrating a flow equation. Leading divergences are then extracted from the two-particle interaction vertex by solving the linearized gap equation. In the employed truncated-unity formulation, the interaction vertices are expressed in terms of the lattice harmonics [30, 31, 32, 29, 33, 34] giving us precise access to the explicit forms of the leading instabilities. We calculated up to possible forth-nearest-neighbor pairings, with a transfer momentum discretization of $108^{2}$ momentum points. In addition, each transfer momentum point of the Green’s function has been obtained by averaging over $51^{2}$ points of an even finer mesh centered around each of the coarser momentum grid points [33].
@@
Given the model construction, we are interested in weak coupling, $U\ll W$, with $W=6t$ the bandwidth. In practice, we operate at $21/3\leq U/W\leq 1/$ 2 in order to obtain numerically tractable instabilities; for the same reason we focus on band fillings in the vicinity of the van-Hove singularity located $n_{\rm vH}=5/4$ where effects of weak interactions are strongest.
@@
In Fig. 2 (a) we present the TUFRG phase diagram for the AFM Hubbard–Kitaev model. We find a spin-density wave (SDW) at and near van-Hove filling, surrounded at lower and higher fillings by a two-fold degenerate superconducting instability with $d$-wave pairing symmetry. The dominant superconducting pairing is between nearest neighbors, with a relatively small next-nearest neighbor contribution and negligible pairings on third- and forth-nearest neighbors. As discussed below, this pairing corresponds to the chiral $d$-wave state well-known from other hexagonal-lattice systems with $E_{2g}$ irrep. Its real-space structure is shown in Fig. 3(b), while the momentum-space order parameters of the different spin and orbital states are plotted in the Supplementary Material (SM). The results for $U/t=2,2.5,3$ are qualitatively similar, Fig. 2(a); stronger interactions slightly increase the filling range of the magnetic phase, but mostly increase the critical scale $\Lambda_{c}$ and thus stabilize the symmetry-broken states. Importantly, our numerical results are stable over an extended $U$ range and do not require any fine-tuning.
In Fig. 2(b) we present the analogous phase diagram for the FM case. We again find a SDW at and near van-Hove filling, surrounded at lower and higher fillings by a now three-fold degenerate superconducting instability, which corresponds to $p$-wave pairing, which, in the same way as in the AFM case, is primarily weighted on the nearest-neighbor bonds. We emphasize that this $p$-wave state is different from the standard chiral $p+ip$ state with $E_{1u}$ irrep known from the $D_{6h}$ point group, details are discussed below.
@@
For $U/t=2$ and fillings $n\leq 0.6$, the FRG does not find any instability down to the lowest cutoff considered, $\Lambda_{\text{min}}=10^{-6}$. A finer momentum discretization would be needed to determine whether an instability occurs at even smaller scales. However, given that $\Lambda_{\text{min}}$ is already small, the practical conclusion is that there is no instability down to lowest temperatures.
@@
The SDW phases in both phase diagrams correspond to commensurate (incommensurate) spin order at (near) van-Hove filling, respectively; further details are relegated to the SM.
@@
Mean-field theory of fractionalized superconductivity.— In order to study the properties of the emergent SC state, we treat the Hubbard–Kitaev model using BCS-type mean-field theory. TUFRG delivers an effective bilinear interaction corresponding to the superconducting instability, which we can directly input into a Bogoliubov–de Gennes (BdG) Hamiltonian. To this end, we use the standard definition of the superconducting order parameter [35],
@@
where the set of matrices $\{\sigma_{0},\bm{\sigma}\}$ act on the usual spin subspace, with the identity matrix $\sigma_{0}$ and the Pauli matrices $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$. All the spatial (and therefore sublattice) pairing information is contained in the coefficients; $\Psi$ is the spin-singlet pairing and $\bm{d}=(d_{x},d_{y},d_{z})$ are the three spin-triplet pairings. The representation theory of the effective model reveals that the corresponding total group is $O_{h}\times\mathbb{Z}_{2}$; for details see SM.
AFM case.— The previously mentioned two-fold degenerate $d$-wave states form a set of basis states of the $E_{2g}$ irrep of the group $O_{h}\times\mathbb{Z}_{2}$. We can write the superconducting term of the mean-field Hamiltonian, as known from the standard Hubbard model [36, 37, 34], as
@@
Through a calculation of the Berry curvature using numerical integration [38], we find Chern number $\mathcal{C}=4$ for fillings of the superconducting regime both below and above the SDW phase. $\mathcal{C}=4$ matches the results in the literature for a chiral $d$-wave superconductor with dominant nearest-neighbor pairing [39, 34]. In the SM, we also show examples of corresponding ribbon spectra with four chiral in-gap states per edge, as expected from bulk–boundary correspondence.
Superconducting order parameters are conventionally named after their angular-momentum winding around the $\Gamma$ point ($s,p,\ldots$ wave). In systems that have quantum indices with non-trivial spatial structure, such as the honeycomb lattice with two distinct sublattices, this must be done with care. We can use the angular-momentum label consistently, as long as we consider only those components of the Cooper pairings that map to themselves when the sublattice indices are swapped. On the honeycomb lattice, these pairings are the symmetric intra-sublattice pairing $\Delta\sim a^{\dagger}a^{\dagger}+b^{\dagger}b^{\dagger}$, and the symmetric inter-sublattice pairing $\Delta\sim a^{\dagger}b^{\dagger}+b^{\dagger}a^{\dagger}$. This allows us to identify this superconductor as a chiral $d$-wave, for further information see SM.
FM case.—The previously mentioned three-fold degenerate superconducting instability leads to a spin-triplet $p$-wave state on the nearest-neighbor bonds. These three states span the $T_{1u}$ irrep of the total group. The pairing part of the mean-field Hamiltonian can be written as
@@
where $m\in\{x,y,z\}$ for each of the three degenerate SC states. For instance, when $\vec{\Delta}_{z}\propto(1,-1,0)$, the pairing amplitude $\Delta^{3}_{z}=0$ across the Kitaev $z$ bond, Eq. (6), as depicted in Fig. 3 (c). The other two degenerate states are obtained by permuting the entries of $\vec{\Delta}$, fixing $\Delta_{m}^{\alpha}=0$ for $m=\alpha$. In the same way we classified the AFM instability as $d$-wave above, we can classify this SC order parameter as $p$-wave.
In order to find the stable superconducting state, we perform a Landau-Ginzburg analysis [40, 36] (see SM for details) which identifies the so-called chiral superposition [41, 42, 43], $\vec{\Delta}_{x}+\varepsilon\vec{\Delta}_{y}+\varepsilon^{*}\vec{\Delta}_{z}$, with $\varepsilon=e^{i2\pi/3}$ as energetically most favorable. The resulting state is fully gapped. By performing the analogous topological analysis, we find $\mathcal{C}=4$ in the superconducting phase for fillings below the SDW phase. For the phase above the SDW phase, the Chern number is $\mathcal{C}=2$. In the SM, we show examples of corresponding ribbon spectra.
@@
Given the results of Refs. 17, 18, the SC phase established in this paper will transit into a non-fractionalized superconducting phase upon increasing $J_{\rm K}$. The $T=0$ SC–SC transition is a confinement transition; it can be either of first order or continuous. In the latter case, the simplest theory is that of the condensation of a bosonic field corresponding to Kondo screening [3, 4]. Given the $\mathbb{Z}_{2}$ nature of the spin liquid’s gauge field, this quantum phase transition is of Ising type [15]. Since the normal-state Fermi surface reconstructs at the parent FL–FL transition, the fermiology of the two superconducting states (Chern numbers etc.) can be expected to be different. Interestingly, the $\mathbb{Z}_{2}$ confinement transition is only well defined at $T=0$ in two space dimensions, such that the continuous SC–SC transition becomes a crossover between two quasi-long-range-ordered states at finite $T$. A more detailed study of this transition is left for future work; some of its signatures have been considered in Ref. 15.
@@
Summary.— We have shown that the Kitaev–Kondo lattice admits a fractionalized superconducting phase SC, where BCS-type superconductivity of conduction electrons coexists with a spin-liquid background and an emergent gauge structure. This phase is separated from a more conventional (heavy-fermion) superconductor by a confinement transition driven by Kondo screening (or the breakdown thereof); this transition is hence the superconducting version of the FL–FL transition discussed in the context of Kondo-breakdown quantum criticality. The pairing of SC is driven by the magnetic excitations of the Kitaev spin liquid, i.e., by Majorana glue. Our FRG analysis has revealed the pairing symmetry to be topologically non-trivial and of chiral spin-singlet (chiral spin-triplet) type for AFM (FM) Kitaev interactions, for band fillings away from but near van-Hove filling. Our arguments for an SC phase straightforwardly extend to other Kitaev spin liquids; in three-dimensional systems [44] where true superconductivity exists at finite $T$ the $\mathbb{Z}_{2}$ confinement transition can be expected to persist at finite $T$ as well.
@@
More broadly, we have established a new pairing mechanism – the fluctuations of a fractionalized spin liquid – leading to topological superconductivity. Future theory work should establish such pairing beyond the perturbative limit. Experimentally, studying hetero-systems comprised of frustrated local moments and conduction electrons appears most promising to identify the phases proposed here.
@@
Note added.— While completing this work, we became aware of Ref. 45 which also studies the emergence of weak-coupling superconductivity in the FM Kitaev–Kondo lattice. Their results are broadly consistent with ours, however, they only consider small dopings near half filling, related to the phase we find at fillings below the van-Hove singularity, with Chern number $2$ instead of 4. A very different type of fractionalized superconductor appears in Ref. 46: This has fractional charges and spinor order, but no deconfined gauge field.

### v2

# Fractionalized Superconductivity Mediated by Majorana Fermions in the Kitaev-Kondo Lattice

@@ Abstract
Fractionalized phases of matter in space dimensions $d\geq 2$ attract widespread interest in condensed matter research: They are characterized by emergent gauge fields, topological order, and fractionalized excitations which constitute novel quasiparticles [1, 2]. Fractionalization is often discussed for bulk insulators such as gapped spin liquids and fractional Quantum-Hall states. Significantly less work has been devoted to fractionalized metals and superconductors, mainly because the presence of bulk charge carriers renders both conceptual and experimental distinctions between fractionalized and conventional phases more subtle than in the simpler insulating cases.
One class of fractionalized metals whose existence can be rigorously established theoretically is that of fractionalized Fermi liquids (FL) [3, 4]: A two-band model where electrons in one strongly correlated band form a fractionalized spin liquid of local moments, e.g. because of frustration, and are weakly exchange-coupled to a second band of weakly correlated electrons realizes an FL phase. The FL phase is characterized by the coexistence of fractionalized spin excitations, and their associated gauge field, as well as charge carriers with charge $|e|$ and spin $1/2$ whose Fermi volume violates Luttinger’s theorem [5]. A stronger coupling between the two bands leads to Kondo screening of the local moments, resulting in a more conventional heavy Fermi-liquid (FL) phase. Consequently, the transition between FL and FL is an instance of a Kondo-breakdown quantum phase transition [6, 7, 3, 8, 9], accompanied by a jump in the Fermi volume but no spontaneous symmetry breaking.
The focus of this paper is on constructing a superconducting cousin of the FL phase, i.e., a fractionalized superconductor (SC), where superconductivity coexists with topological order. In fact, this has been hypothesized early on to exist in underdoped cuprate superconductors [10]. More recently, it has been discussed [11] in scenarios of the cuprate pseudogap representing a fractionalized Fermi liquid [12, 13, 14]. A transition between fractionalized and conventional superconductivity (SC) has also been proposed to occur in an iron-based superconductor [15].
@@
In this paper we discuss and solve a tractable model of fractionalized superconductivity. We consider the Kitaev–Kondo lattice model, Fig. 1, constructed from the honeycomb-lattice spin-liquid model of Kitaev [16] and a band of conduction electrons. This model has been shown to display a FL–FL transition which – at the mean-field level – is masked by non-trivial but non-fractionalized superconductivity at intermediate Kondo coupling [17, 18]. Here we focus on the fractionalized weak-coupling regime where we can perturbatively derive an effective model for conduction-electron interactions mediated by the Majorana fermions of the spin liquid. We solve this model using the functional renormalization group (FRG) technique [19, 20] and find stable superconductivity of either chiral $d$-wave ($p$-wave) type for ferromagnetic (antiferromagnetic) Kitaev coupling, respectively, over a range of band fillings. These pairing states coexist with the fractionalized spin liquid in the local-moment sector, rendering this SC phase a natural descendant of the model’s FL phase.
@@
in standard notation, with $S_{i}^{\alpha}$ representing the Kitaev spin at site $i$ and $s^{\alpha}_{i}=c_{i\sigma}^{\dagger}\tau^{\alpha}_{\sigma\sigma^{\prime}}c_{i\sigma^{\prime}}/2$ the conduction-electron spin operator with Pauli matrices $\tau^{\alpha}$. The first term represents the conduction-electron kinetic energy, the second the Kitaev coupling among the spin-$1/2$ local moments, with $\langle ij\rangle_{\alpha}$ an $\alpha$ bond on the lattice ($\alpha=x,y,z$), and the last the local Kondo coupling. A chemical potential $\mu$ is applied to the conduction electrons to control their filling, $n_{c}=\frac{1}{2N}\sum_{i\sigma}c_{i\sigma}^{\dagger}c_{i\sigma}$, where $N$ is the number of unit cells. Half filling corresponds to $n_{c}=1$. We will concentrate on $n_{c}\geq 1$; phases for $n_{c}\leq 1$ are related by particle–hole symmetry.
@@
The Kitaev model $\mathcal{H}_{K}$ alone describes an exactly solvable $\mathbb{Z}_{2}$ spin liquid [16]. Its degrees of freedom are itinerant “matter” Majorana fermions and static $\mathbb{Z}_{2}$ gauge fluxes. Here we focus on isotropic Kitaev couplings, $K^{\alpha}\equiv K$, for which the matter-Majorana spectrum is gapless, and on isotropic Kondo couplings, $J_{\rm K}^{\alpha}\equiv J_{\rm K}$. Upon adding the Kondo coupling, the $\mathbb{Z}_{2}$ fluxes are no longer conserved, and the solubility is spoiled.
We quickly summarize the results of Refs. 17, 18 where the Kitaev–Kondo lattice model $\mathcal{H}$ has been studied using perturbative and mean-field techniques. The model hosts a $\mathbb{Z}_{2}$ fractionalized Fermi liquid [3] for $J_{\rm K}\ll K,t$ because the Kitaev spin liquid is stable against a small coupling to conduction electrons. Conversely, the model realizes a conventional heavy Fermi liquid for $K\ll J_{\rm K}\sim t$ (or $K\ll T_{\rm K}$ where $T_{\rm K}$ is the Kondo temperature) with Kondo screening of the local moments. According to the mean-field results, the transition between these two symmetric metallic phases is masked by superconductivity, in accordance with general considerations [3]. At the mean-field level, superconductivity sets in concomitantly with Kondo screening, hence the superconducting phase is non-fractionalized (but may have non-trivial topological band invariants). The precise character of the superconducting phases depends on the mean-field scheme and differs between Refs. 17 and 18. We note that the Kitaev–Kondo lattice has also been utilized to describe heterostructures composed of graphene and $\alpha$-RuCl3 [25].
@@
Figure 2: FRG phase diagrams of the (a) AFM and (b) FM Hubbard–Kitaev model as a function of critical scale $\Lambda_{c}$ vs. filling $n_{c}$. A magnetically ordered phase (SDW) at and near van-Hove filling $n_{\rm vH}=5/4$ is surrounded on both sides by a chiral $d$-wave, and $p$-wave superconductor, respectively. Shown are results for $U/t=2$, 2.5 and 3. The crosses indicate the absence of any divergence, i.e., the Fermi liquid prevails down to scales of $\Lambda=10^{-6}$.
@@
FRG analysis of effective model.— We analyze the leading instabilities of the Hubbard–Kitaev model by virtue of the truncated-unity FRG (TUFRG) [19, 20] method, using the recently published FRG codebase divERGe [29]. FRG interpolates between the bare Hubbard–Kitaev interaction and a low-energy two-particle interaction vertex, by means of iteratively integrating a flow equation. Leading divergences are then extracted from the two-particle interaction vertex by solving the linearized gap equation. In the employed truncated-unity formulation, the interaction vertices are expressed in terms of the lattice harmonics [30, 31, 32, 29, 33, 34] giving us precise access to the explicit forms of the leading instabilities. We calculated up to possible forth-nearest-neighbor pairings, with a transfer momentum discretization of $108^{2}$ momentum points. A further $51^{2}$ refinement is used to evaluate the Green’s functions [33].
@@
Given the model construction, we are interested in weak coupling, $U\ll W$, with bandwidth $W=6t$. In practice, we operate at $21/3\leq U/W\leq 1/$ 2 in order to obtain numerically tractable instabilities; for the same reason we focus on band fillings in the vicinity of the van-Hove singularity located $n_{\rm vH}=5/4$ where effects of weak interactions are strongest.
@@
In Fig. 2 we present the TUFRG phase diagrams for the (a) AFM and (b) FM Hubbard-Kitaev models. We find a spin-density wave (SDW) at and near van-Hove filling, surrounded at lower and higher fillings by superconducting phases. The results for $U/t=2,2.5,3$ are qualitatively similar, Fig. 2; stronger interactions slightly increase the filling range of the magnetic phase, but mostly increase the critical scale $\Lambda_{c}$ and thus stabilize the symmetry-broken states. Importantly, our numerical results are stable over an extended $U$ range and do not require any fine-tuning.
@@
For the FM Hubbard-Kitaev model with $U/t=2$ and fillings $n_{c}\leq 0.6$, the FRG does not find any instability down to the lowest cutoff considered, $\Lambda_{\text{min}}=10^{-6}$. A finer momentum discretization would be needed to determine whether an instability occurs at even smaller scales.
@@
The SDW phases in both phase diagrams correspond to commensurate (incommensurate) spin order at (near) van-Hove filling, respectively; further details are relegated to the SM [35].
@@
Mean-field theory of fractionalized superconductivity.— In order to study the properties of the emergent SC state, we treat the Hubbard–Kitaev model using BCS-type mean-field theory. TUFRG delivers an effective bilinear interaction corresponding to the superconducting instability, which we can directly input into a Bogoliubov–de Gennes (BdG) Hamiltonian. We define the superconducting order parameter [36],
@@
where the identity matrix $\sigma_{0}$ and the Pauli matrices $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ act on the spin subspace. All the spatial (and therefore sublattice) pairing information is contained in the coefficients; $\Psi$ is the spin-singlet pairing and $\bm{d}=(d_{x},d_{y},d_{z})$ are the three spin-triplet pairings. The representation theory of the effective model reveals that the corresponding total group is $O_{h}\times\mathbb{Z}_{2}$; for details see SM.
AFM case.— The SC state is composed of two-fold degenerate $d$-wave states, which form a set of basis states of the $E_{2g}$ irrep of the group $O_{h}\times\mathbb{Z}_{2}$. We can write the superconducting term of the mean-field Hamiltonian, as known from the standard Hubbard model [37, 38, 34], as
@@
The dominant superconducting pairing is between nearest neighbors, with a relatively small next-nearest neighbor contribution and negligible pairings on third- and forth-nearest neighbors. This pairing corresponds to the chiral $d$-wave state well-known from other hexagonal-lattice systems with $E_{2g}$ irrep [39, 37, 34]. Its real-space structure is shown in Fig. 3(b), while the momentum-space order parameters of the different spin and orbital states are plotted in the Supplementary Material (SM) [35]. The SM also contains the rationale behind the $d$-wave naming convention for the honeycomb lattice.
@@
Through a calculation of the Berry curvature using numerical integration [40], we find Chern number $\mathcal{C}=4$ for both superconducting phases. $\mathcal{C}=4$ matches the results in the literature for a chiral $d$-wave superconductor with dominant nearest-neighbor pairing [41, 34]. In the SM, we also show examples of corresponding ribbon spectra with four chiral in-gap states per edge, as expected from bulk–boundary correspondence.
FM case.—The three-fold degenerate superconducting instability leads to a spin-triplet $p$-wave state primarily on the nearest-neighbor bonds. These three states span the $T_{1u}$ irrep of the total group. The pairing part of the mean-field Hamiltonian can be written as
@@
where $m\in\{x,y,z\}$ for each of the three degenerate SC states. For instance, when $\vec{\Delta}_{z}\propto(1,-1,0)$, the pairing amplitude $\Delta^{3}_{z}=0$ across the Kitaev $z$ bond, Eq. (6), as depicted in Fig. 3 (c). The other two degenerate states are obtained by permuting the entries of $\vec{\Delta}$, fixing $\Delta_{m}^{\alpha}=0$ for $m=\alpha$. In the same way we classified the AFM instability as $d$-wave above, we can classify this SC order parameter as $p$-wave, see the SM for further details. We emphasize that this $p$-wave state is different from the standard chiral $p+ip$ state with $E_{1u}$ irrep known from the $D_{6h}$ point group.
In order to find the stable superconducting state, we perform a Landau-Ginzburg analysis [42, 37] (see SM [35] for details) which identifies the so-called chiral superposition [43, 44, 45], $\vec{\Delta}_{x}+\varepsilon\vec{\Delta}_{y}+\varepsilon^{*}\vec{\Delta}_{z}$, with $\varepsilon=e^{i2\pi/3}$ as energetically most favorable. The resulting state is fully gapped. By performing the analogous topological analysis, we find $\mathcal{C}=4$ in the superconducting phase for fillings below the SDW phase. For the phase above the SDW phase, the Chern number is $\mathcal{C}=2$. In the SM, we show examples of corresponding ribbon spectra.
@@
Given the results of Refs. 17, 18, the SC phase established in this paper will transit into a non-fractionalized superconducting phase upon increasing $J_{\rm K}$. The $T=0$ SC–SC transition is a confinement transition; it can be either of first order or continuous. In the latter case, the simplest theory is that of the condensation of a bosonic field corresponding to Kondo screening [3, 4]. Given the $\mathbb{Z}_{2}$ nature of the spin liquid’s gauge field, this quantum phase transition is of Ising type [15]. Since the normal-state Fermi surface reconstructs at the parent FL–FL transition, the fermiology of the two superconducting states (Chern numbers etc.) can be expected to be different. Interestingly, the $\mathbb{Z}_{2}$ confinement transition is only well defined at $T=0$ in two space dimensions, such that the continuous SC–SC transition becomes a crossover between two quasi-long-range-ordered states at finite $T$. A detailed study of this transition is left for future work; some of its signatures have been considered in Ref. 15.
@@
Summary.— We have shown that the Kitaev–Kondo lattice admits a fractionalized superconducting phase SC, where BCS-type superconductivity of conduction electrons coexists with a spin-liquid background and an emergent gauge structure. This phase is separated from a more conventional (heavy-fermion) superconductor by a confinement transition driven by Kondo screening (or its breakdown); this transition is hence the superconducting version of the FL–FL transition discussed in the context of Kondo-breakdown quantum criticality. The pairing of SC is driven by the magnetic excitations of the Kitaev spin liquid, i.e., by Majorana glue. Our FRG analysis has revealed the pairing symmetry to be topologically non-trivial and of chiral spin-singlet (chiral spin-triplet) type for AFM (FM) Kitaev interactions, for band fillings away from but near van-Hove filling. Our arguments for an SC phase straightforwardly extend to other Kitaev spin liquids; in three-dimensional systems [46] where true superconductivity exists at finite $T$ the $\mathbb{Z}_{2}$ confinement transition can be expected to persist at finite $T$ as well.
@@
More broadly, we have established a new pairing mechanism – the fluctuations of a fractionalized spin liquid – leading to topological superconductivity. Future work should establish such pairing beyond the perturbative limit and in non-Kitaev systems. Experimentally, studying hetero-systems comprised of frustrated local moments and conduction electrons appears most promising to identify the phases proposed here.
@@
Note added.— While completing this work, we became aware of Ref. 47 which also studies weak-coupling superconductivity in the FM Kitaev–Kondo lattice. Their results are broadly consistent with ours, however, they only consider small dopings near half filling, related to the phase we find at fillings below the van-Hove singularity, with Chern number $2$ instead of 4. A very different type of fractionalized superconductor appears in Ref. 48: This paper investigates a Kondo lattice based on the Yao-Lee model where the gauge field is deconfined across the entire phase diagram by construction.